OpenStudy (anonymous):

ABCD-rhombus. Angle PAQ = 56 degrees. Find angle B.

OpenStudy (anonymous):

a figure would help

OpenStudy (anonymous):

It looks like triangle ABP has two right angles?

OpenStudy (anonymous):

No, I just painted angle B

OpenStudy (anonymous):

 do its \[\angle B= 90\]

OpenStudy (anonymous):

\[B \neq90\]

OpenStudy (campbell_st):

look at the quadrilateral APCQ angle QAP = 56 angle AQC = CPA = 90 then by angle sum of a quadrilateral angle PCQ = 124 next look then by the property of opposite angles in a rhombus are equal angle BAD = 124 angles DAQ + BAP = 124 - 56 = 68

OpenStudy (anonymous):

Thank you campbell_st:)

OpenStudy (campbell_st):

using opposite angles of a rhombus are equal and angle BAD = angle BCD = 124 then angles ABC + ADC = 360 - 248 = 112 2ABC = 112 ABC = 56

OpenStudy (anonymous):

thanks again))